# Retroreflective surface optimisation for optical cavities with custom mirror profiles

### v1

@@ II Mode parameterisation and optimisation
The paraxial equation (Eq. 2) is satisfied by Gaussian beam solutions, which are particularly convenient for Fabry-Pérot cavities because they are transversely localised and satisfy the reflection boundary conditions imposed by spherical mirrors. We will be concerned with radially symmetric mirrors, and will therefore use the Laguerre-Gauss modes, which may be written
@@
where $\lambda=2\pi/k$ is the wavelength, $L^{\alpha}_{n}$ are the associated Laguerre polynomials, and $z_{0}$ is the Rayleigh range of the beam. The coordinate origin has been fixed to lie at the centre of the smallest waist of the beam. The Laguerre-Gauss modes are indexed by radial index $l\in\mathbb{N}_{0}$ and azimuthal index $m\in\mathbb{Z}$. They form a complete basis ($\left\{\epsilon_{l,m}\right\}$}[37], which means that any field satisfying the paraxial equation can be written as a linear superposition of these basis modes,
$\epsilon^{\pm}(\bm{r})=\sum_{s=0}^{S}C_{s}\epsilon^{\pm}_{s}(\bm{r}),$ (5)
provided the basis size $S+1$ is sufficiently large, where $(l,m)$ is mapped, for notational convenience, to a single index $s$, and $C_{s}$ is the complex coefficient of the basis state $\epsilon^{\pm}_{s}$. We can thus identify any solution to the paraxial equation with a complex vector $\vec{C}$
@@
The process by which the coefficients of $\vec{C}$ are optimised depends upon the metric of performance that is being optimised. For the general case where only the performance metric (but not the gradient) can be calculated for a given set of coefficients, one may use local or global optimisation. For a substantial number of interesting cases, the performance metric may be written as the product and quotient of expectation values of Hermitian matrices. In this case, the gradient can be evaluated directly (see App. A), which allows for faster local optimisation.
[…]
@@ III Obtaining mirror topography from retroreflection
To construct the mirror surface, we need to find the equiphase contour of the target electric field. The phase of the electric field is composed of the phase of the paraxial mode $\epsilon_{\mathrm{T}}$, and the propagating wave phase $\exp(\mp ikz)$ (see Eq. 1). We will take the mirror at axial coordinate $z_{m}>0$ as an example, and find the radially-dependent protrusion $f(r)$, which, by convention, points towards the centre of the cavity. For each coordinate $\bm{r}=(r,\phi,z_{m})$ on the mirror plane, the incoming paraxial mode $\epsilon^{+}_{T}$ has a phase, whose differences across the mirror profile should be offset by the running wave phase incident on the varying surface protrusion. Mathematically,
$kf(r)\mod 2\pi=-\mathrm{arg}\left[\epsilon^{+}_{T}(r,z_{m})\right].$ (7)
@@
The mirror surface on the other side of the cavity will have the same protrusion (also pointing towards the centre of the cavity) due to the enforced axial reflection symmetry of the mode. Strictly speaking, $f(r)$ is only an equiphase contour on the assumption that the phase of $\epsilon_{\mathrm{T}}$ changes only negligibly as $z$ changes between the nominal mirror plane and the protruded surface. We do not correct this assumption because the discrepancy generally has little impact, and because the mode mixing method, which we use to calculate the cavity eigenmodes, also makes this assumption [39].
@@
There is, however, a complication with the logic of retroreflective optimisation. The target mode $\epsilon_{\mathrm{T}}$ is an eigenmode on the assumption that it can be retroreflected. However, if there is any clipping loss, the chosen mode cannot be perfectly retroreflected because the field leaking beyond the mirror cannot be retroreflected. This means that we must calculate the true eigenmode of the cavity, $\epsilon_{\mathrm{E}}$, instead of assuming that $\epsilon_{\mathrm{T}}$ is an eigenmode. Additionally, the performance calculated for $\epsilon_{\mathrm{E}}$ might not be strictly the maximum possible value, only a lower bound to it, because it was $\epsilon_{\mathrm{T}}$, not $\epsilon_{\mathrm{E}}$, that was optimised. In practice, most sensible metrics of cavity performance will discourage clipping loss that significantly reduces the number of cavity round trips. This means that the clipping loss of the target mode $\epsilon_{\mathrm{T}}$ is likely to be low, so little power in $\epsilon_{\mathrm{T}}$ will leak beyond the mirrors, and therefore $\epsilon_{\mathrm{E}}$ is likely to be very similar to $\epsilon_{\mathrm{T}}$. Empirical evidence for this is presented in App. B.
@@
To calculate the non-Gaussian eigenmodes of cavities with non-spherical mirrors, we use mode mixing, as described in [40], which has been used in a variety of theoretical [17, 38, 19, 41, 42] and experimental studies [43, 44, 30]. This requires the evaluation of the surface through Eq. 7 at multiple radial coordinates. The modulus is taken to give the smoothest surface (smallest vertical change between adjacent points). In our calculations, if the sum of any two adjacent changes exceeded a phase of $\pi$, the number of samples was deemed insufficient, and the process repeated with double the sampling density.
@@ IV.1 Identifying the performance metric
We begin with the simple example of a single emitter in the centre of an optical cavity, with our goal to maximise the probability of extracting a single photon from the emitter in a single attempt. This figure of merit is very important for emitter-cavity systems used as single photon sources [45], but also for systems that produce emitter-photon entanglement to use in quantum repeaters and quantum networks [7], where the operation rate depends strongly on the photon emission probability [46, 8, 9, 47, 48].
@@
$C_{\mathrm{int}}=\frac{g^{2}}{\kappa_{\mathrm{int}}\gamma},$ (8)
@@
$P_{\mathrm{ext}}=\frac{\sqrt{2C_{\mathrm{int}}-1}}{\sqrt{2C_{\mathrm{int}}+1}}$ (9)
where this limit is satisfied in infinite time [49]. In practice, This bound holds well provided the photon is generated over a time-scale longer than the ‘critical time’ $t_{\mathrm{crit}}=\max(1/\kappa,\kappa/g^{2})$ [50, 51], which is satisfied in many experimental situations [52].
An alternative expression for $C_{\mathrm{int}}$ emphasises the role of the mode profile through the ‘inverse effective area’, $A^{-1}_{\mathrm{eff}}(\bm{r_{e}})={\epsilon^{+}}^{2}\left(\bm{r_{e}}\right)/\int_{\Pi_{z_{e}}}{\epsilon^{+}}^{2}\left(\bm{r}\right)\,d^{2}\bm{r}$, where $\Pi_{z_{e}}$ is the transverse plane containing the emitter, which represents the ratio of the intensity of the forward-running cavity mode at the position $\bm{r_{e}}$ of the emitter to the total forward-running power. The expression is
@@
where $\lambda$ is the emitter transition wavelength, and $\mathcal{L}_{\mathrm{int}}$ is the round-trip power loss to non-transmissive sources (see [53, 54] for a derivation, where the internal loss $\mathcal{L}_{\mathrm{int}}$, rather than total loss, has been used in order to express $C_{\mathrm{int}}$ rather than the cooperativity $C$, and appropriate quantities have been substituted to match the current notation). Note that the expression assumes, as we shall do throughout, that the emitter is placed perfectly at the antinode of the axial standing wave, which will require a maximum of $\lambda/4$ axial translation of the emitter from its nominal position. It also assumes that the branching ratio of the excited state emission to the cavity mode, which accounts for both the spontaneous emission branching ratio, and the polarisation overlap with the cavity mode, is unity. In the case that the branching ratio is not unity, the right hand side of Eq. 10 should be multiplied by that branching ratio.
@@ IV.2 Evaluating the performance metric
To use the retroreflective optimisation method, we need to be able to calculate $C_{\mathrm{int}}$ directly from the cavity mode, and, in order to use more efficient optimisation algorithms, we can calculate the gradient of $C_{\mathrm{int}}$ as well. To do this, we can write the inverse effective area at the emitter $A^{-1}_{\mathrm{eff}}(\bm{r_{e}})$ and internal loss $\mathcal{L}_{\mathrm{int}}$ as (unnormalised) matrices.
The inverse area matrix elements are given
$\hat{A^{-1}_{\mathrm{eff}}}_{s^{\prime},s}(\bm{r}_{e})=\left(\epsilon^{+}_{s^{\prime}}(\bm{r_{e}})\right)^{*}\epsilon^{+}_{s}(\bm{r_{e}}),$ (11)
@@
$\left(\hat{\mathcal{L}}_{\mathrm{int}}\right)_{s^{\prime},s}$ $=2\left[\hat{\mathcal{I}}-\left(1-\frac{1}{2}\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}\right)\hat{\rho}^{{\dagger}}\hat{\rho}\right]_{s^{\prime},s},$ (12)
$\hat{\rho}_{s^{\prime},s}$ $=\int_{\Pi_{m}}\left(\epsilon^{+}_{s^{\prime}}(\bm{r})\right)^{*}\epsilon^{+}_{s}(\bm{r})\,d\Pi,$ (13)
where $\hat{\rho}$ is the clipping matrix of the aperture defined by the edge of one mirror with $\Pi_{m}$ the transverse plane of that mirror, $\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}$ is the internal round trip loss when clipping from outside the mirror apertures is not counted, and $\hat{\mathcal{I}}$ is the identity matrix. For the spherical apertures considered in this manuscript, the elements of $\hat{\rho}$ can be evaluated using the techniques of Vajente [55]. The details of the matrix derivations are given in App. C.
@@ IV.3 Possible improvements in performance
We first consider the case that the mirrors are spherical, and thus the cavity modes are Gaussian. Given a cavity length $L$, there is a minimum viable diameter $D_{\mathrm{crit}}$, below which every possible Gaussian mode sustains clipping losses that exceed the other internal losses of the cavity $\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}$. In this case, the cavity performance is low. We can derive an approximate expression for $D_{\mathrm{crit}}$
$D_{\mathrm{crit}}=2\sqrt{\frac{\lambda L}{\pi}}\sqrt{-\frac{1}{2}\log\left(\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}\right)},$ (14)
from the power clipping approximation [28] (see App. D). If the diameter exceeds $D_{\mathrm{crit}}$, then $C_{\mathrm{int}}$ may be increased by reducing the mirror radius of curvature, and thus focusing the mode more tightly on the emitter, until the increasingly-divergent mode sustains clipping losses on the mirrors that do become significant when compared to $\mathcal{L}_{\mathrm{int}}^{\mathrm{clip}\rightarrow 0}$ [60]. The spherical mirror shapes, and consequent Gaussian fundamental mode, mean that only the divergence of the cavity mode can be optimised (which determines the waist of the central focus and the waist on the mirrors); there is no flexibility to further control the transverse profile of the mode as it always has a Gaussian decay.
@@
We investigate the potential of retroreflective optimisation to improve the internal cooperativity compared to any cavity with spherical mirrors, and compared to cavities designed through the ansatz method. In Fig. 1, we study an example situation, and perform the optimisation over a range of cavity geometries to explore trends in the optimisation performance. To perform retroreflective optimisation for a given cavity geometry, we first found the best radius of curvature to use for spherical mirrors, which sets the Laguerre-Gauss basis modes of the optimisation. We then constructed the appropriate matrices to calculate $C_{\mathrm{int}}$ (see Sec. IV) and optimised the real parameter vector $\vec{C}$ to find $\epsilon_{\mathrm{T}}$ with an L-BFGS algorithm [61]. We used the first 171 $m=0$ Laguerre-Gauss modes to construct the matrices, with the first 136 of these corresponding to free coefficients of $\vec{C}$ (the coefficients of the remaining basis states were set to zero). The number of free coefficients was chosen to be less than the basis size to ensure that the clipping loss of the highest occupied basis state of $\epsilon_{\mathrm{T}}$ (corresponding to the final free parameter) was correctly calculated by accounting for its overlap with basis states of somewhat larger $l$-index. The basis size and number of free parameters were chosen to ensure convergence of optimised $C_{\mathrm{int}}$ over the range of geometries investigated in Fig. 1.
@@
Moving beyond the single example, we can perform the same optimisation procedure for a range of geometries, here varying the length and the mirror diameter. In Fig. 1e), we show the performance achievable with spherical mirrors and with optimised mirrors. The most obvious features are the regions of high and low performance, roughly delineated by the white line depicting $D_{\mathrm{crit}}$, despite its crude derivation that ignores mode-mixing effects. Retroreflective optimisation can subtly move the exact boundary between the regions, but, similar to the case with spherical mirrors, cannot generally find good performance if $D<D_{\mathrm{crit}}$. However, retroreflective optimisation can significantly increase $C_{\mathrm{int}}$ in the high performance region ($D>D_{\mathrm{crit}}$). The scale of improvement tends to increase with increasing mirror diameter before saturating at an improvement factor of roughly 3, which is broadly constant across a large range of length scales. This suggests that ideal mirror shaping could increase internal cooperativity by roughly half an order of magnitude in the central emitter scenario provided the mirror is large enough. This increase is considerable, but the optimised mirror shapes are highly complex, and therefore likely not ideal designs for practical manufacture.

### v2

@@ II Mode parameterisation and optimisation
The paraxial equation (Eq. (2)) is satisfied by Gaussian beam solutions, which are particularly convenient for Fabry-Pérot cavities because they are transversely localised and satisfy the reflection boundary conditions imposed by spherical mirrors. We will be largely concerned with radially symmetric mirrors, and will therefore use the Laguerre-Gauss modes, which may be written
@@
where $\lambda=2\pi/k$ is the wavelength, $L^{\alpha}_{n}$ are the associated Laguerre polynomials, and $z_{0}$ is the Rayleigh range of the beam. The coordinate origin has been fixed to lie at the centre of the smallest waist of the beam. The Laguerre-Gauss modes are indexed by radial index $l\in\mathbb{N}_{0}$ and azimuthal index $m\in\mathbb{Z}$.
We will also optimise cavities in one transverse dimension (i.e. suppressing functional dependence in, and derivatives of $y$). For these cases, the paraxial equation can be solved with Hermite-Gauss modes
$\epsilon^{HG,\pm}_{n}\left(\bm{r}\right)=\sqrt{\sqrt{\frac{2}{\pi}}\frac{1}{w(z)2^{n}n!}}H_{n}\left(\frac{\sqrt{2}x}{w(z)}\right)\exp\left\{-\frac{x^{2}}{w^{2}(z)}-i\left(\pm\frac{kx^{2}}{2R_{u}(z)}\mp\left(n+\frac{1}{2}\right)\Phi_{G}(z)\right)\right\},$ (5)
where $H_{n}$ are the physicist’s Hermite polynomials and $n\in\mathbb{N}_{0}$ is the transverse index of the mode. Both Laguerre-Gauss and Hermite Gauss states form complete basis sets (Laguerre-Gauss over a 2D transverse plane, Hermite-Gauss modes over a single transverse dimension) [37]. This means that any field satisfying the paraxial equation can be written as a linear superposition of the basis modes in the appropriate domain,
$\epsilon^{\pm}(\bm{r})=\sum_{s=0}^{S}C_{s}\epsilon^{\Lambda,\pm}_{s}(\bm{r}),$ (6)
provided the basis size $S+1$ is sufficiently large, where $\Lambda$ is either $LG$ or $HG$, the mode indices $(l,m)$ or $n$ are mapped to a single index $s$, and $C_{s}$ is the complex coefficient of the basis state $\epsilon^{\Lambda,\pm}_{s}$. We can thus identify any solution to the paraxial equation with a complex vector $\vec{C}$
@@
The process by which the coefficients of $\vec{C}$ are optimised depends upon the metric of performance that is being optimised. For the general case where only the performance metric (but not the gradient) can be calculated for a given set of coefficients, one may use local or global optimisation. For a substantial number of interesting cases, the performance metric may be written as the product and quotient of expectation values of Hermitian matrices. In this case, the gradient can be evaluated directly (see App. B), which allows for faster local optimisation.
@@ III Obtaining mirror topography from retroreflection
To construct the mirror surface, we need to find the equiphase contour of the target electric field. The phase of the electric field is composed of the phase of the paraxial mode $\epsilon_{\mathrm{T}}$, and the propagating wave phase $\exp(\mp ikz)$ (see Eq. (1)). For each coordinate $\bm{r}=(\bm{r}_{\perp},z_{m_{\pm}})$ on the plane for the positive (negative) $z$ mirror, where $\bm{r}_{\perp}$ is some parameterisation of the transverse coordinate, the incoming paraxial mode $\epsilon^{+}_{T}$ has a phase, whose differences across the mirror profile should be offset by the running wave phase incident on the varying surface protrusion $f_{\pm}(\bm{r}_{\perp})$ which, by convention, points towards the centre of the cavity. Mathematically,
$kf_{\pm}(\bm{r}_{\perp})\mod 2\pi=\mp\mathrm{arg}\left[\epsilon^{+}_{T}(r,z_{m_{\pm}})\right].$ (8)
@@
Strictly speaking, $f_{\pm}(\bm{r}_{\perp})$ is only an equiphase contour on the assumption that the phase of $\epsilon_{\mathrm{T}}$ changes only negligibly as $z$ changes between the nominal mirror plane and the protruded surface. We do not correct this assumption because the discrepancy generally has little impact, and because the mode mixing method, which we use to calculate the cavity eigenmodes, also makes this assumption [39].
@@
There is, however, a complication with the logic of retroreflective optimisation. The target mode $\epsilon_{\mathrm{T}}$ is an eigenmode on the assumption that it can be retroreflected. However, if there is any clipping loss, the chosen mode cannot be perfectly retroreflected because the field leaking beyond the mirror cannot be retroreflected. This means that we must calculate the true eigenmode of the cavity, $\epsilon_{\mathrm{E}}$, instead of assuming that $\epsilon_{\mathrm{T}}$ is an eigenmode. Additionally, the performance calculated for $\epsilon_{\mathrm{E}}$ might not be strictly the maximum possible value, only a lower bound to it, because it was $\epsilon_{\mathrm{T}}$, not $\epsilon_{\mathrm{E}}$, that was optimised. In practice, most sensible metrics of cavity performance will discourage clipping loss that significantly reduces the number of cavity round trips. This means that the clipping loss of the target mode $\epsilon_{\mathrm{T}}$ is likely to be low, so little power in $\epsilon_{\mathrm{T}}$ will leak beyond the mirrors, and therefore $\epsilon_{\mathrm{E}}$ is likely to be very similar to $\epsilon_{\mathrm{T}}$. Empirical evidence for this is presented in App. C.
Finally, the surface construction method can encounter problems for arbitrary modes with two transverse dimensions, because these can feature phase vortices, which are points where the amplitude goes to zero, but the phase change encountered on a path encircling the vortex does not sum to zero (instead summing to a multiple of $2\pi$). The equiphase surface of a mode with a vortex can require infinitely sharp drops of size $\lambda/2$ in order to maintain phase continuity, often unavoidably through regions of significant mode intensity. While such a mirror surface is a mathematical possibility, it is not a practical solution, and the infinite gradient cliffs would violate the paraxial approximation on which the method rests. To avoid the problem of phase vortices, we restrict our optimisations to situations where the phase only varies with one transverse parameter. For the majority of cases, this parameter will be the cylindrical polar radius $r$ with the assumption of complete azimuthal rotation symmetry, which is guaranteed by using only Laguerre-Gauss modes of $m=0$. We will also use the Hermite-Gauss modes, whose transverse parameter is the Cartesian $x$ coordinate, for cases that break radial symmetry but maintain translational symmetry along the other transverse direction ($y$). Though this is a restriction, we note that many cavities are close to satisfying one of these cases, and for those that are not, one may still obtain qualitative or pedagogical insight.
@@
To calculate the non-Gaussian eigenmodes of cavities with non-spherical mirrors, we use mode mixing, as described in [40], which has been used in a variety of theoretical [17, 38, 19, 41, 42] and experimental studies [43, 44, 30]. This requires the evaluation of the surface through Eq. (8) at multiple transverse coordinates. The modulus is taken to give the smoothest surface (smallest vertical change between adjacent points). In our calculations, if the sum of any two adjacent changes exceeded a phase of $\pi$, the number of samples was deemed insufficient, and the process repeated with double the sampling density.
@@ IV.1 Identifying the performance metric
We begin with the simple example of a single emitter in the centre of an axially symmetric optical cavity, with our goal to maximise the probability of extracting a single photon from the emitter in a single attempt. This figure of merit is very important for emitter-cavity systems used as single photon sources [45], but also for systems that produce emitter-photon entanglement to use in quantum repeaters and quantum networks [7], where the operation rate depends strongly on the photon emission probability [46, 8, 9, 47, 48].
@@
$C_{\mathrm{int}}=\frac{g^{2}}{2\kappa_{\mathrm{int}}\gamma},$ (9)
@@
$P_{\mathrm{ext}}=\frac{\sqrt{2C_{\mathrm{int}}+1}-1}{\sqrt{2C_{\mathrm{int}}+1}+1}$ (10)
where this limit is satisfied in infinite time [49]. In practice, This bound holds well provided the photon is generated over a time-scale longer than the ‘critical time’ $t_{\mathrm{crit}}=\max(1/\kappa,\kappa/g^{2})$ [50, 51], which is satisfied in many experimental situations [52]. For a graphical example of how $C_{\mathrm{int}}$ affects $P_{\mathrm{ext}}$, see App. A.
An alternative expression for $C_{\mathrm{int}}$ emphasises the role of the mode profile through the ‘inverse effective area’, $A^{-1}_{\mathrm{eff}}(\bm{r_{e}})={\epsilon^{+}}^{2}\left(\bm{r_{e}}\right)/\int_{P_{z_{e}}}{\epsilon^{+}}^{2}\left(\bm{r}\right)\,d^{2}\bm{r}$, where $P_{z_{e}}$ is the transverse plane containing the emitter, which represents the ratio of the intensity of the forward-running cavity mode at the position $\bm{r_{e}}$ of the emitter to the total forward-running power. The expression is
@@
where $\lambda$ is the emitter transition wavelength, and $\mathcal{L}_{\mathrm{int}}$ is the round-trip power loss to non-transmissive sources (see [53, 54] for a derivation, where the internal loss $\mathcal{L}_{\mathrm{int}}$, rather than total loss, has been used in order to express $C_{\mathrm{int}}$ rather than the cooperativity $C$, and appropriate quantities have been substituted to match the current notation). Note that the expression assumes, as we shall do throughout, that the emitter is placed perfectly at the antinode of the axial standing wave, which will require a maximum of $\lambda/4$ axial translation of the emitter from its nominal position. It also assumes that the branching ratio of the excited state emission to the cavity mode, which accounts for both the spontaneous emission branching ratio, and the polarisation overlap with the cavity mode, is unity. In the case that the branching ratio is not unity, the right hand side of Eq. (11) should be multiplied by that branching ratio.
For the case of a cavity with one transverse dimension (specifically, a cavity extended in the $y$-direction so that the $y$-dependence of the cavity mode is negligible), $C_{\mathrm{int}}$ cannot be directly calculated. Instead, we can use the cartesian separability of the mode in this case to assert $A^{-1}_{\mathrm{eff}}(\bm{r_{e}})=L^{-1}_{\mathrm{eff}}(\bm{r_{e}})L^{-1}_{\mathrm{y,eff}}$, where the inverse effective length in the $x$-direction $L^{-1}_{\mathrm{eff}}(\bm{r_{e}})={\epsilon^{+}}^{2}\left(\bm{r_{e}}\right)/\int_{X_{z_{e}}}{\epsilon^{+}}^{2}\left(\bm{r}\right)\,d\bm{r}$, where $X_{z_{e}}$ is the $x$-axis in the transverse plane containing the emitter. For a cavity that is infinitely extended in $y$, the inverse effective length in the y direction, $L^{-1}_{\mathrm{y,eff}}$, is zero, but for any real cavity, $L^{-1}_{\mathrm{y,eff}}$ will be finite. We can then optimise the one-transverse dimension cavity using the metric
$C_{\mathrm{int}}/L^{-1}_{\mathrm{y,eff}}=\frac{3\lambda^{2}L^{-1}_{\mathrm{eff}}(\bm{r_{e}})}{\pi\mathcal{L}_{\mathrm{int}}}.$ (12)
This means that $L^{-1}_{\mathrm{y,eff}}$ never need be calculated to perform optimisation in cases where the $y$-dependence is negligible, but for a real cavity with non-infinite $y$ extent, the value of $L^{-1}_{\mathrm{y,eff}}$ could be supplied to yield $C_{\mathrm{int}}$.
@@ IV.2 Evaluating the performance metric
To use the retroreflective optimisation method, we need to be able to calculate $C_{\mathrm{int}}$ or $C_{\mathrm{int}}/L^{-1}_{\mathrm{y,eff}}$ directly from the cavity mode, and, in order to use more efficient optimisation algorithms, we can calculate their gradients as well. To do this, we can write the inverse effective area (length) at the emitter $A^{-1}_{\mathrm{eff}}(\bm{r_{e}})$ ($L^{-1}_{\mathrm{eff}}(\bm{r_{e}})$) and internal loss $\mathcal{L}_{\mathrm{int}}$ as (unnormalised) matrices.
The inverse area and inverse length matrix elements are given
$\hat{A^{-1}_{\mathrm{eff}}}_{s^{\prime},s}(\bm{r}_{e})=\left(\epsilon^{LG,+}_{s^{\prime}}(\bm{r_{e}})\right)^{*}\epsilon^{LG,+}_{s}(\bm{r_{e}}),\quad\hat{L^{-1}_{\mathrm{eff}}}_{s^{\prime},s}(\bm{r}_{e})=\left(\epsilon^{HG,+}_{s^{\prime}}(\bm{r_{e}})\right)^{*}\epsilon^{HG,+}_{s}(\bm{r_{e}}),$ (13)
@@
$\hat{\mathcal{L}}_{\mathrm{int}}$ $=\hat{\mathcal{L}}_{\mathrm{int}}^{+}+\hat{\mathcal{L}}_{\mathrm{int}}^{-}$ (14)
$\hat{\mathcal{L}}_{\mathrm{int}}^{\pm}$ $=\hat{\mathcal{I}}-\left(1-\frac{1}{2}\mathcal{L}_{\mathrm{abs}}\right)\left(\hat{\rho}^{\pm}\right)^{{\dagger}}\hat{\rho^{\pm}},$
$\hat{\rho}^{\pm}_{s^{\prime},s}$ $=\int_{\Pi_{m_{\pm}}}\left(\epsilon^{\Lambda,+}_{s^{\prime}}(\bm{r})\right)^{*}\epsilon^{\Lambda,+}_{s}(\bm{r})\,d\Pi,$
where $\hat{\mathcal{L}}_{\mathrm{int}}^{\pm}$ is the internal loss matrix of the mirror at positive (negative) $z$ coordinate, $\hat{\mathcal{I}}$ is the identity matrix, $\mathcal{L}_{\mathrm{abs}}$ is the round trip absorption loss (including scattering), which is the component of the internal round trip loss that does not include clipping, $\hat{\rho}^{\pm}$ is the clipping matrix of the aperture defined by the edge of the mirror at positive (negative) $z$, and $\Pi_{m_{\pm}}$ is the finite domain (a surface for simulations with two transverse dimensions, and a length for a single transverse dimension) of the mirror at positive (negative) $z$. For the spherical apertures considered in the radially symmetric simulations, the elements of $\hat{\rho}^{\pm}$ can be evaluated using the techniques of Vajente [55]. The details of the matrix derivations are given in App. D.
@@ IV.3 Possible improvements in performance
We first consider the case that the mirrors are spherical, and thus the cavity modes are Gaussian. Given a cavity length $L$, there is a minimum viable diameter $D_{\mathrm{crit}}$, below which every possible Gaussian mode sustains clipping losses that exceed the other internal losses of the cavity $\mathcal{L}_{\mathrm{abs}}$. In this case, the cavity performance is low. We can derive an approximate expression for $D_{\mathrm{crit}}$
$D_{\mathrm{crit}}=2\sqrt{\frac{\lambda L}{\pi}}\sqrt{-\frac{1}{2}\log\left(\mathcal{L}_{\mathrm{abs}}\right)},$ (15)
from the power clipping approximation [28] (see App. E). If the diameter exceeds $D_{\mathrm{crit}}$, then $C_{\mathrm{int}}$ may be increased by reducing the mirror radius of curvature, and thus focusing the mode more tightly on the emitter, until the increasingly-divergent mode sustains clipping losses on the mirrors that do become significant when compared to $\mathcal{L}_{\mathrm{abs}}$ [60]. The spherical mirror shapes, and consequent Gaussian fundamental mode, mean that only the divergence of the cavity mode can be optimised (which determines the waist of the central focus and the waist on the mirrors); there is no flexibility to further control the transverse profile of the mode as it always has a Gaussian decay.
@@
We investigate the potential of retroreflective optimisation to improve the internal cooperativity compared to any cavity with spherical mirrors, and compared to cavities designed through the ansatz method. In Fig. 1, we study an example situation, and perform the optimisation over a range of cavity geometries to explore trends in the optimisation performance. To perform retroreflective optimisation for a given cavity geometry, we first found the best radius of curvature to use for spherical mirrors, which sets the Laguerre-Gauss basis modes of the optimisation. We then constructed the appropriate matrices to calculate $C_{\mathrm{int}}$ (see Sec. IV) and optimised the real parameter vector $\vec{C}$ to find $\epsilon_{\mathrm{T}}$ with an L-BFGS algorithm [61]. We used the first 171 $m=0$ Laguerre-Gauss modes to construct the matrices, with the first 136 of these corresponding to free coefficients of $\vec{C}$ (the coefficients of the remaining basis states were set to zero). The number of free coefficients was chosen to be less than the basis size to ensure that the clipping loss of the highest occupied basis state of $\epsilon_{\mathrm{T}}$ (corresponding to the final free parameter) was correctly calculated by accounting for its overlap with basis states of somewhat larger $l$-index. The basis size and number of free parameters were chosen to ensure convergence of optimised $C_{\mathrm{int}}$ over the range of geometries investigated in Fig. 1. Details of the basis sizes required for convergence and the parameters used for all simulations in this manuscript can be found in App. F.
@@
Moving beyond the single example, we can perform the same optimisation procedure for a range of geometries, here varying the length and the mirror diameter. In Fig. 1e), we show the performance achievable with spherical mirrors and with optimised mirrors. The most obvious features are the regions of high and low performance, roughly delineated by the white line depicting $D_{\mathrm{crit}}$, despite its crude derivation that ignores mode-mixing effects. Retroreflective optimisation can subtly move the exact boundary between the regions, but, similar to the case with spherical mirrors, cannot generally find good performance if $D<D_{\mathrm{crit}}$. However, retroreflective optimisation can significantly increase $C_{\mathrm{int}}$ in the high performance region ($D>D_{\mathrm{crit}}$). The scale of improvement tends to increase with increasing mirror diameter before saturating at an improvement factor of roughly 3, which is broadly constant across a large range of length scales. This suggests that ideal mirror shaping could increase internal cooperativity by roughly half an order of magnitude in the central emitter scenario provided the mirror is large enough. This increase is considerable, but the optimised mirror shapes are highly complex, and therefore likely not ideal designs for practical manufacture (see Sec. VI for further discussion on this point).
Finally, we note that optimisation also changes the $g$ and $\kappa$ parameters of the cavity-QED system. Most directly, the more intense focussing of the cavity mode on the emitter increases the emitter-cavity coupling rate $g$. The optimised cavity $\kappa_{\mathrm{int}}$ typically remains close to the minimum limit set by $\mathcal{L}_{\mathrm{abs}}$, but the increased $C_{\mathrm{int}}$ means the optimum transmission of the outcoupling mirror will increase, increasing $\kappa$ [49]. While the impacts of changes to $g$ and $\kappa$ on the photon extraction probability are encapsulated by Eqs. (9) and (10), these rates are also important for achieving strong coupling ($g>\kappa,\gamma$), which is important for observing vacuum-Rabi oscillations, as well as affecting the resonant selectivity of the cavity [62], and its sensitivity to birefringence [63]. If the experimental application were to place constraints on the acceptable values of $g$ and $\kappa$ beyond a desire to maximise $P_{\mathrm{ext}}$, these constraints could be included in the metric optimised to select $\epsilon_{\mathrm{T}}$.